OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 3y + 12 -3y + 12 3y - 12 -3y - 12

OpenStudy (dan815):

You have the left side expression and the right side expression equal to each other For example like A=B then if you add or subtract or divide or multiply both sides by the same quantity, they should still be the same to each other Based of this principal we have -x - 3y = -12 add x to both sides -x+x-3y=-12+x -3y=-12+x add 12 to both sides -3y+12=-12+12+x -3y+12=x